# Sparse Autoencoder Neural Operators: Model Recovery in Function Spaces

### v2

@@ Appendix B Appendix - Additional Background
Notations We denote scalars as non-bold, lower-case a, vectors as bold-lower-case $\bm{a}$}, and matrices as upper-case letters $\bm{A}$}.
@@
Sparse model recovery The bi-level optimization (Def.˜1.2) takes a general form. When the inner level objective $\mathcal{L}({\bm{x}},{\bm{v}},\theta)=\frac{1}{2}\|{\bm{x}}-{\bm{D}}{\bm{v}}\|_{2}^{2}+\mathcal{R}({\bm{v}})$, this problem is reduced to the classical alternating-minimization dictionary learning [42, 43]. In this classical form, sparse coding is widely applied in science and engineering; e.g., in signal processing [56], image denoising [57], and extracting interpretable gene modules [58, 59]. When the model parameters $\bm{D}$} are known, the problem is reduced to recovering the sparse representation by solving $\min_{{\bm{z}}\in\mathbb{R}^{p}}\ \tfrac{1}{2}\|{\bm{x}}-{\bm{D}}{\bm{z}}\|_{2}^{2}+\lambda\mathcal{R}({\bm{z}})$, where $\mathcal{R}({\bm{z}})$ is a sparse regularizer; setting $\mathcal{R}({\bm{z}})=\|{\bm{z}}\|_{1}$, the optimization problem is reduced down to lasso [60], or referred to as basis pursuit [61], solving via proximal gradient descent [62] such as iterative shrinkage-thresholding algorithm (ISTA) [63] and its fast momentum-based version [64].
Sparse ReLU autoencoders For shallow ReLU autoencoders [26], the formulation reduces to
@@
This connection between bilevel optimization and neural networks, highlighted in [44, 55], has been explored extensively in the context of sparse autoencoders. Prior works have studied the gradient dynamics of variants of ReLU networks when the encoder is shallow [49, 50, 52] or deep, as in learned ISTA architectures [48]. Others have analyzed model recovery when the encoder is deep and iterative with hard-thresholding nonlinearity (JumpReLU [27][54].
Overall, optimizing or recovering sparse generative linear models has attracted sustained attention from two main deep learning communities over the past years; a) Unrolling learning: using optimization formulations to design and theoretically study neural architectures [46, 47, 48, 49, 50, 51, 52, 53, 54], and b) Sparse interpretability or mechanistic interpretability: leveraging sparse models to interpret and analyze the internal representations of complex larger networks [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34]. We briefly expand on both works.
Unrolled learning The early connections between sparse coding and deep learning trace back to sparse energy-based deep models [65, 66] and the pioneering work of LISTA [45] in constructing sparsifying recurrent neural networks. This line of work, known as unfolding [67] or unrolling [68], designs neural network layers based on iterations of an algorithm that solves an optimization problem. This connection has enabled researchers to use optimization models as a proxy to theoretically study accelerated convergence [51, 46], gradient dynamics [50, 52, 49, 53], and model recovery in both shallow [50] and deep neural networks [54, 48]. Furthermore, several works have explored the theoretical connections between convolutional neural networks and convolutional sparse coding [69, 70], or more generally, between deeply sparse signal representations and deep neural networks [71]. An example of a deep unrolled JumpReLU neural network for sparse model recovery can be formulated as the following bilevel optimization problem.
@@
for $t=1,\ldots,T$, where ${\bm{z}}_{0}=\bm{0}$, $\alpha$ is the step size, and $\text{JumpReLU}_{\lambda}({\bm{z}})={\bm{z}}\cdot\mathbbm{1}_{{\bm{z}}>\lambda}$, where $\mathbbm{1}$ is the indicator function, and $\lambda$ controls the sparsity level of the representation. The inner mapping is now a deep/iterative encoder, which for simplicity we denote by ${\bm{z}}_{T}=f_{\theta}({\bm{x}})$. The outer objective enforces the structure of the decoder and the loss function. The bilevel optimization in eq.˜5 can be mapped to the following neural network autoencoder architecture, which uses a recurrent and residual encoder. The synthetic experiments in this paper use this encoder in its architecture for sparse model recovery.
@@
Sparse Interpretability Interpretability in deep learning [72], particularly through extracting concepts learned by deep neural networks [21], has gained prominence in the era of large models. Motivated by the linear representation hypothesis (LRH) [24, 28], sparse autoencoders (SAEs) are widely used to recover interpretable structures from internal representations of large language models [25, 26, 27, 31] and vision models [32]. Recent studies show that the architecture of SAEs imposes inductive biases on which concepts are recoverable [55]. Building on this, we investigate conditions under which neural networks and neural operators recover similar concepts, and we explore their respective inductive biases. Our formulation of sparse model recovery in function spaces enables the use of SAEs operating on functions to interpret large neural operators.
Convolutional Setting Sparse generative models can be extended to sparse convolutional generative models (Def.˜B.1), where the concepts are locally and sparsely appearing in the data.
Definition B.1 (Sparse Convolutional Generative Models).
@@
  • Kernel Integration: For $t=0,\ldots,T-1$, this is a non-local integral kernel operator that maps representation at one layer $\{{\bm{v}}_{t}:D_{t}\rightarrow\mathbb{R}^{d_{v_{t}}}\}$ to the next $\{{\bm{v}}_{t+1}:D_{t+1}\rightarrow\mathbb{R}^{d_{v_{t+1}}}\}$ (see Def.˜B.2).
@@
Definition B.2 (Kernel integral operator $\mathcal{K}$ (restated from [13, 14])).
@@
Definition B.3 (Fourier transform).
Let $\mathcal{F}$ be the Fourier transform of the function ${\bm{v}}:D\rightarrow\mathbb{R}^{d_{v}}$, whose inverse is denoted by $\mathcal{F}^{-1}$ on the function ${\bm{w}}:D\rightarrow\mathbb{C}^{d_{w}}$. We have
@@
Appendix C Appendix - Additional Experimental Results
C.1 Experimental Setup
@@
Sparse Generative Model  The results shown in Fig.˜6 are from a synthetic dataset that follows a sparse generative model (Def.˜1.1). The atoms of the ground-truth dictionary ${\bm{D}}^{*}\in\mathbb{R}^{1000\times 1500}$ are drawn from a standard normal distribution and then $\ell_{2}$-normalized. The dataset consists of 50,00050{,}000 samples ${\bm{x}}\in\mathbb{R}^{1000}$. Each latent code ${\bm{z}}\in\mathbb{R}^{1500}$ has a total sparsity of $20$. The amplitudes of the non-zero elements are drawn from a sub-Gaussian distribution $\mathcal{N}(15,1)$.
@@
SAE  For SAE (SAE-MLP) and L-SAE (L-SAE-MLP) models, the encoder implements $f_{\theta}({\bm{x}})$ following Def.˜1.2 using a deep unrolled JumpReLU network (eqs.˜5 and 6) with $T=50$ layers, as described in the sparse model recovery framework (Def.˜1.2). The non-linearity is Hard-Thresholding with a threshold of $\lambda=0.5$, and the algorithm’s internal step-size is $\alpha=0.2$. The decoder reconstructs data as $\hat{{\bm{x}}}={\bm{D}}{\bm{z}}$ following Def.˜1.3. The dictionary weights for SAE and L-SAE are initialized with a noisy ($\sigma=0.02)$ version of the ground-truth dictionary. For the L-SAE, the 1000-dimensional input is lifted to a 1200-dimensional space. Both models are trained to minimize the Mean Squared Error (MSE) loss using the ADAM optimizer with a learning rate of $\eta=10^{-3}$.
@@
Sparse Convolutional Generative Model  The results shown in Fig.˜2 are from a synthetic dataset that follows a sparse convolutional generative model (Def.˜B.1). The ground-truth dictionary $\bm{D}$$^{*}$ consists of $p=5$ kernels with spatial support $h=99$. Each kernel is a multi-channel signal with 64 input channels. The values for each kernel are drawn from a standard normal distribution and then normalized. The dataset consists of 50,000 samples ${\bm{x}}\in\mathbb{R}^{64\times 1000}$. Each sample is generated by convolving a single (i.e., sparsity of 1) randomly chosen kernel from the dictionary with a sparse feature map, whose non-zero amplitudes are drawn from a sub-Gaussian distribution.
@@
SAE-CNN  For SAE-CNN and L-SAE-CNN models, the encoder implements the convolutional version of the deep unrolled network (eq.˜7) with $T=50$ layers, as described in the sparse model recovery framework (Def.˜1.2). The non-linearity is Hard-Thresholding with a threshold of $\lambda=10$, and the algorithm’s internal step-size is $\alpha=0.01$. The dictionary for both SAE-CNN and L-SAE-CNN are initialized with a noisy ($\sigma=0.05$) version of the ground-truth kernels. For L-SAE-CNN, the 64-channel input is lifted to a 128-dimensional space. Both models are trained to minimize the Mean Squared Error (MSE) loss using the SGD optimizer with a learning rate of $\eta=0.04$.
@@
SAE-FNO For SAE-FNO and L-SAE-FNO models, the encoder implements a deep unrolled network with $T=50$ layers, as described in the sparse functional model recovery framework (Def.˜2.2). The non-linearity is Hard-Thresholding with a threshold of $\lambda=10$, and the algorithm’s internal step-size is $\alpha=0.01$. The dictionary kernels are initialized with a noisy ($\sigma=0.05)$ version of the ground-truth. For L-SAE-FNO, the 64-channel input is lifted to a 128-dimensional space. Both SAE-FNO and L-SAE-FNO are trained to minimize the Mean Squared Error (MSE) loss using the SGD optimizer with a learning rate of $\eta_{L}=20.04$. All experiments are based on these settings, except Fig.˜2b and Fig.˜3.
Fig.˜2b uses a modified experimental setup from the sparse convolutional generative model described above. The data consists of single channel signal with 1-sparse codes ($p=1$ kernel), where the amplitudes of non-zero elements are drawn from a sub-Gaussian distribution $\mathcal{N}(5,1)$. Training uses threshold $\lambda=0.2$, step size $\alpha=0.05$, and ADAM optimizer with learning rates $2\times 10^{-5}$ (SAE-CNN) and $0.01$ (SAE-FNO).
Fig.˜3 uses the experimental setup of Sparse Convolutional Generative Model with $1\times$, $2\times$, $4\times$, and $8\times$ upsampling rates. The upsampling process is as follows: original low-resolution data ($m=1000$) is first upsampled by inserting zeros at intervals corresponding to the upsampling rate (e.g., for $4\times$ upsampling, 3 zeros are inserted between every original sample, creating a signal of length 4000). This process introduces high-frequency spectral replicas of the original signal. The expanded signal is then processed through a 10th-order Butterworth low-pass filter using zero-phase filtering (filtfilt) to remove aliasing and ensure smooth interpolation.
C.2 Experimental Results
Additional experimental results are shown in Figs.˜4, 5, 6 and 7.
Fig. 4: Lifting as a preconditioner. Lifting accelerates learning.
Fig. 5: Lifting. When the lifting operator satisfies the orthogonal condition ${\bm{L}}^{\top}{\bm{L}}={\bm{I}}$, the lifted SAE-CNN (L-SAE-CNN) exhibits equivalent learning dynamics to the SAE-CNN .
(a) Dictionary Recovery Error
@@ C.2 Experimental Results
Fig. 6: SAEs vs. L-SAEs. Learning the lifting operator $\bm{L}$} accelerates model recovery. (a) Dictionary recovery error converges faster for L-SAE-MLP, confirming the preconditioning effect of lifting; (b) Reconstruction loss follows similar convergence trends as dictionary recovery error; (c) Lifting encourages the effective dictionary $\bm{D}$} to learn more orthogonal (less correlated) atoms early in training, creating a more isotropic loss landscape that accelerates recovery; (d) The lifting operator $\bm{L}$} initially becomes less orthogonal, allowing the dictionary $\bm{D}$} to become more orthogonal.
(a) Without Upsampling.
(b) $2\times$ Upsampling.
(c) $4\times$ Upsampling.
(d) $8\times$ Upsampling.
[…]
Appendix D Appendix - Theoretical Analysis

### v3

@@ Appendix A Appendix - Additional Background
Notations We denote scalars as non-bold-lower-case a, vectors as bold-lower-case $\bm{a}$}, and matrices as upper-case letters $\bm{A}$}.
Sparse model recovery has attracted interest from two representation learning communities: a) unrolled learning, using optimization-inspired architectures for learning sparse representation (Ablin et al., 2019; Malézieux et al., 2022; Tolooshams and Ba, 2022; Nguyen et al., 2019; Arora et al., 2015; Chen et al., 2018; Rangamani et al., 2018; Tolooshams et al., 2020; Rambhatla et al., 2018); and b) sparse interpretability or mechanistic interpretability, employing SAEs to extract concepts from large models (Elhage et al., 2022; Huben et al., 2023; Bricken et al., 2023b; Rajamanoharan et al., 2024; Park et al., 2024; Templeton et al., 2024; Lieberum et al., 2024; Gao et al., 2025; Fel et al., 2025; Marks et al., 2025; Karvonen et al., 2025).
@@
Sparse model recovery The bi-level optimization (Definition 2.2) takes a general form. When the inner level objective $\mathcal{L}({\bm{x}},{\bm{v}},\theta)=\frac{1}{2}\|{\bm{x}}-{\bm{D}}{\bm{v}}\|_{2}^{2}+\mathcal{R}({\bm{v}})$, this problem is reduced to the classical alternating-minimization dictionary learning (Agarwal et al., 2016; Chatterji and Bartlett, 2017). In this classical form, sparse coding is widely applied in science and engineering; e.g., in signal processing (Elad, 2010), image denoising (Aharon et al., 2006), and extracting interpretable gene modules (Cleary et al., 2017, 2021). When the model parameters $\bm{D}$} are known, the problem is reduced to recovering the sparse representation by solving $\min_{{\bm{z}}\in\mathbb{R}^{p}}\ \tfrac{1}{2}\|{\bm{x}}-{\bm{D}}{\bm{z}}\|_{2}^{2}+\lambda\mathcal{R}({\bm{z}})$, where $\mathcal{R}({\bm{z}})$ is a sparse regularizer; setting $\mathcal{R}({\bm{z}})=\|{\bm{z}}\|_{1}$, the optimization problem is reduced down to lasso (Tibshirani, 1996), or referred to as basis pursuit (Chen et al., 2001), solving via proximal gradient descent (Parikh and Boyd, 2014) such as iterative shrinkage-thresholding algorithm (ISTA) (Daubechies et al., 2004) and its fast momentum-based version (Beck and Teboulle, 2009).
Sparse ReLU autoencoders For shallow ReLU autoencoders (Bricken et al., 2023b), the formulation reduces to
@@
This connection between bilevel optimization and neural networks, highlighted in (Tolooshams, 2023; Hindupur et al., 2025), has been explored extensively in the context of sparse autoencoders. Prior works have studied the gradient dynamics of variants of ReLU networks when the encoder is shallow (Nguyen et al., 2019; Arora et al., 2015; Rangamani et al., 2018) or deep, as in learned ISTA architectures (Tolooshams and Ba, 2022). Others have analyzed model recovery when the encoder is deep and iterative with hard-thresholding nonlinearity (JumpReLU (Rajamanoharan et al., 2024)(Rambhatla et al., 2018).
Overall, optimizing or recovering sparse generative linear models has attracted sustained attention from two main deep learning communities over the past years; a) Unrolling learning: using optimization formulations to design and theoretically study neural architectures (Ablin et al., 2019; Malézieux et al., 2022; Tolooshams and Ba, 2022; Nguyen et al., 2019; Arora et al., 2015; Chen et al., 2018; Rangamani et al., 2018; Tolooshams et al., 2020; Rambhatla et al., 2018), and b) Sparse interpretability or mechanistic interpretability: leveraging sparse models to interpret and analyze the internal representations of complex larger networks (Elhage et al., 2022; Huben et al., 2023; Bricken et al., 2023b; Rajamanoharan et al., 2024; Park et al., 2024; Templeton et al., 2024; Lieberum et al., 2024; Gao et al., 2025; Fel et al., 2025; Marks et al., 2025; Karvonen et al., 2025). We briefly expand on both works.
Unrolled learning The early connections between sparse coding and deep learning trace back to sparse energy-based deep models (Ranzato et al., 2007, 2008) and the pioneering work of LISTA (Gregor and LeCun, 2010) in constructing sparsifying recurrent neural networks. This line of work, known as unfolding (Hershey et al., 2014) or unrolling (Monga et al., 2021), designs neural network layers based on iterations of an algorithm that solves an optimization problem. This connection has enabled researchers to use optimization models as a proxy to theoretically study accelerated convergence (Chen et al., 2018; Ablin et al., 2019), gradient dynamics (Arora et al., 2015; Rangamani et al., 2018; Nguyen et al., 2019; Tolooshams et al., 2020), and model recovery in both shallow (Arora et al., 2015) and deep neural networks (Rambhatla et al., 2018; Tolooshams and Ba, 2022). Furthermore, several works have explored the theoretical connections between convolutional neural networks and convolutional sparse coding (Papyan et al., 2017a, b), or more generally, between deeply sparse signal representations and deep neural networks (Ba, 2020). An example of a deep unrolled JumpReLU neural network for sparse model recovery can be formulated as the following bilevel optimization problem.
@@
for $t=1,\ldots,T$, where ${\bm{z}}_{0}=\bm{0}$, $\alpha$ is the step size, and $\text{JumpReLU}_{\lambda}({\bm{z}})={\bm{z}}\cdot\mathbbm{1}_{{\bm{z}}>\lambda}$, where $\mathbbm{1}$ is the indicator function, and $\lambda$ controls the sparsity level of the representation. The inner mapping is now a deep/iterative encoder, which for simplicity we denote by ${\bm{z}}_{T}=f_{\theta}({\bm{x}})$. The outer objective enforces the structure of the decoder and the loss function. The bilevel optimization in Equation 7 can be mapped to the following neural network autoencoder architecture, which uses a recurrent and residual encoder. The synthetic experiments in this paper use this encoder in its architecture for sparse model recovery.
@@
Convolutional Setting Sparse generative models can be extended to sparse convolutional generative models (Definition A.1), where the concepts are locally and sparsely appearing in the data.
Definition A.1 (Sparse Convolutional Generative Models).
@@
  • Kernel Integration: For $t=0,\ldots,T-1$, this is a non-local integral kernel operator that maps representation at one layer $\{{\bm{v}}_{t}:D_{t}\rightarrow\mathbb{R}^{d_{v_{t}}}\}$ to the next $\{{\bm{v}}_{t+1}:D_{t+1}\rightarrow\mathbb{R}^{d_{v_{t+1}}}\}$ (see Definition A.2).
@@
Definition A.2 (Kernel integral operator $\mathcal{K}$ (restated from (Kovachki et al., 2023))).
@@
Definition A.3 (Fourier integral operator $\mathcal{K}$ (restated from (Kovachki et al., 2023))).
Define the Fourier integral operator by
$(\mathcal{K}(\phi){\bm{v}}_{t})({\bm{x}})=\mathcal{F}^{-1}\big(R_{\phi}\cdot(\mathcal{F}{\bm{v}}_{t})\big)({\bm{x}}),\qquad\forall{\bm{x}}\in D$ (11)
where $R_{\phi}\coloneqq\mathcal{F}(\kappa)$ is the Fourier transform of a periodic function $\kappa:D\rightarrow\mathbb{C}^{d_{w}\times d_{v}}$ parameterized by $\phi$.
Fourier neural operator (FNO) (Li et al., 2021) models the kernel integral operator with a convolution operator parameterized in Fourier space (Definition A.3, see also Definition A.4 for Fourier transform). The convolution operator parameterizes the kernel $\kappa(x,y)=\kappa(x-y)$ as a complex function $\kappa:D\rightarrow\mathbb{C}^{d_{w}\times d_{v}}$. FNO is used with truncated frequency modes in practice; this has been shown to improve performance and lower sensitivity to change (decreasing) the discretization sampling (Li et al., 2021; Kovachki et al., 2023).
Definition A.4 (Fourier transform).
Let $\mathcal{F}$ the Fourier transform of the function ${\bm{v}}:D\rightarrow\mathbb{R}^{d_{v}}$, whose inverse is denoted by $\mathcal{F}^{-1}$ on the function ${\bm{w}}:D\rightarrow\mathbb{C}^{d_{w}}$. We have
@@
Appendix B Appendix - Additional Experimental Results
B.1 Experimental Setup
@@
Sparse Generative Model  The results shown in Figure 10 are from a synthetic dataset that follows a sparse generative model (Definition 2.1). The atoms of the ground-truth dictionary ${\bm{D}}^{*}\in\mathbb{R}^{1000\times 1500}$ are drawn from a standard normal distribution and then $\ell_{2}$-normalized. The dataset consists of 50,000 samples ${\bm{x}}\in\mathbb{R}^{1000}$. Each latent code ${\bm{z}}\in\mathbb{R}^{1500}$ has a total sparsity of 20. The amplitudes of the non-zero elements are drawn from a sub-Gaussian distribution with a mean of 15 and a standard deviation of 1.0.
@@
SAE  For both SAE (SAE-MLP) and L-SAE (L-SAE-MLP) models, the encoder implements $f_{\theta}({\bm{x}})$ following Definition 2.2 using a deep unrolled JumpReLU network (eq. 7,8) with $T=50$ layers, as described in the sparse model recovery framework (Definition 2.2). The non-linearity is Hard-Thresholding with a threshold of $\lambda=0.5$, and the algorithm’s internal step-size is $\alpha=0.2$. The decoder reconstructs data as $\hat{{\bm{x}}}={\bm{D}}{\bm{z}}$ following Definition 2.3. The dictionary weights for both SAE and L-SAE are initialized with a noisy ($\sigma=0.02)$ version of the ground-truth dictionary. For the L-SAE, the 1000-dimensional input is lifted to a 1200-dimensional space. Both models are trained to minimize the Mean Squared Error (MSE) reconstruction loss using the SGD optimizer with a learning rate of $\eta=10^{-3}$.
@@
Sparse Convolutional Generative Model  The results for convolutional dictionary are from a synthetic dataset that follows a sparse convolutional generative model (Definition A.1). The ground-truth dictionary $\bm{D}$$^{*}$ consists of $p=5$ kernels with spatial support $h=99$. Each kernel is a multi-channel signal with 64 input channels. The values for each kernel are drawn from a standard normal distribution and then normalized. The dataset consists of 50,000 samples ${\bm{x}}\in\mathbb{R}^{64\times 1000}$. Each sample is generated by convolving a single (i.e., sparsity of 1) randomly chosen kernel from the dictionary with a sparse feature map, whose non-zero amplitudes are drawn from a sub-Gaussian distribution.
@@
SAE-CNN  For both SAE-CNN and L-SAE-CNN models, the encoder implements the convolutional version of the a deep unrolled network (Equation 9) with $T=50$ layers, as described in the sparse model recovery framework (Definition 2.2). The non-linearity is Hard-Thresholding with a threshold of $\lambda=10$, and the algorithm’s internal step-size is $\alpha=0.01$. The dictionary kernels for both SAE-CNN and L-SAE-CNN are initialized with a noisy ($\sigma=0.05$) version of the ground-truth kernels. For L-SAE-CNN, the 64-channel input is lifted to a 128-dimensional space. Both models are trained to minimize the Mean Squared Error (MSE) reconstruction loss using the SGD optimizer with a learning rate of $\eta=0.04$.
@@
SAE-FNO For both SAE-FNO and L-SAE-FNO models, the encoder implements a deep unrolled network with $T=50$ layers, as described in the sparse functional model recovery framework (Definition 3.2). The non-linearity is Hard-Thresholding with a threshold of $\lambda=10$, and the algorithm’s internal step-size is $\alpha=0.01$. The dictionary kernels are initialized with a noisy ($\sigma=0.05)$ version of the ground-truth. For L-SAE-FNO, the 64-channel input is lifted to a 128-dimensional space. Both SAE-FNO and L-SAE-FNO are trained to minimize the Mean Squared Error (MSE) reconstruction loss using the SGD optimizer with a learning rate of $\eta_{L}=20.04$.
Figure 9: Lifting. Learning when ${\bm{L}}^{\text{T}}{\bm{L}}={\bm{I}}$.
(a) Dictionary error
@@
Figure 10: SAEs vs. L-SAEs. Learning the lifting operator $\bm{L}$} accelerates model recovery.
Figure 11: SAE-FNO vs. SAE as a function of input discretization. SAE-NO can successfully infer the underlying representations that allow it to reconstruct data at varying resolution beyond the original trained resolution of $28$. The figure shows MSE on the reconstruction.
(a) SAE-FNO (Original rate)
(b) SAE-CNN (Original rate)
(c) SAE-FNO ($2\times$ Upsampling rate)
(d) SAE-CNN ($2\times$ Upsampling rate)
(e) SAE-FNO ($4\times$ Upsampling rate)
(f) SAE-CNN ($4\times$ Upsampling rate)
(g) SAE-FNO ($8\times$ Upsampling rate)
(h) SAE-CNN ($8\times$ Upsampling rate)
Figure 12: SAE-FNO Upsampling Robustness Across Resolutions. We evaluate the ability of SAE-FNO (left column) and SAE-CNN (right column) to infer sparse representations and reconstruct data across varying discretization levels. Each row represents a different upsampling rate ($1\times$ to $8\times$ applied to the input signal). For each panel, the left plot shows the spatial-domain signal reconstruction (red) against the input to the network (blue), and the right plot shows the inference of 1-sparse code supports across 5 kernels. (a-b) Original resolution ($1\times$): Baseline performance at training resolution. Both models accurately reconstruct the signal and recover the correct support. (c-h) Higher upsampling rates ($2\times$, $4\times$, $8\times$): As spatial resolution increases, SAE-FNO maintains accurate reconstruction and code support inference. In contrast, SAE-CNN fails to generalize; its support inference collapses, leading to poor reconstruction. The SAE-FNO’s function space parameterizations enables this generalization across discretizations.
(a) SAE-FNO (modes = 15)
(b) SAE-FNO (modes = 20)
(c) SAE-FNO (modes = 28)
(d) SAE-MLP
Figure 13: Stability of concept characteristics across concept sparsity levels (MNIST). We evaluate the stability of learned concepts of SAE-MLP and SAE-FNO (with spatial sparsity 0.007) by computing the distribution of their Hoyer scores in both the spatial and frequency domains, across varying concept sparsity constraints ($k\in[50,700]$). The y-axis represents the probability density normalized such that the area under the curve equals 1. (a-c) SAE-NO stability: The distribution of both spatial and frequency Hoyer score remains consistent as $k$ varies, regardless of the number of Fourier modes used. (d) SAE-MLP instability: the distributions for SAE-MLP shift drastically with changes in $k$, indicating that the characteristics of learned atoms are highly sensitive to the concept sparsity constraint.
(a) SAE-CNN (kernel size $20\times 20$)
(b) SAE-CNN (kernel size $15\times 15$)
(c) SAE-CNN (kernel size $10\times 10$)
(d) SAE-CNN (kernel size $8\times 8$)
Figure 14: Effect of kernel size on SAE-CNN concept stability (MNIST). We examine Hoyer score densities for SAE-CNN architectures with decreasing kernel sizes. As the kernel size decreases from $20\times 20$ to $8\times 8$, the distributions of atom sparsity become increasingly stable across concept sparsity values. This is expected as restricting the receptive field imposes a stronger inductive bias for locality. Smaller kernels force the network to learn more consistent and localized features.
(a) SAE-FNO (modes = 4)
(b) SAE-FNO (modes = 6)
(c) SAE-FNO (modes = 8)
(d) SAE-MLP
Figure 15: Stability of concept characteristics across concept sparsity levels (CIFAR). CIFAR Hoyer scores for SAE-MLP and SAE-FNO ($s$ = 0.005, 3-sparse)
(a) SAE-FNO (6 sparse)
(b) SAE-FNO (3 sparse)
(c) SAE-FNO (1 sparse)
(d) SAE-MLP
Figure 16: Concept characterization in CIFAR. Hoyer scores for SAE-MLP and SAE-FNO (modes = 6).
(a) $k=50$
(b) $k=100$
(c) $k=200$
(d) $k=500$
(e) $k=700$
(f) $k=1000$
Figure 17: SAE-FNO overcomes orthogonality bias and recovers correlated concepts in MNIST. We compare the distribution of pairwise cosine similarities (correlations) between learned dictionary atoms for SAE-FNO (blue) and SAE-MLP (red) across varying concept sparsity levels ($k$). SAE-FNOs display a broader distribution with heavier tails, indicating a better ability to recover non-orthogonal, correlated concepts.
(a) $k=10$
(b) $k=25$
(c) $k=50$
(d) $k=100$
(e) $k=200$
(f) $k=700$
Figure 18: SAE-FNO overcomes orthogonality bias and recovers correlated concepts in CIFAR. Similar to the MNIST results, we compare the distribution of pairwise cosine similarities between learned dictionary atoms for SAE-FNO (blue) and SAE-MLP (red) across varying concept sparsity levels ($k$). The SAE-FNO models use 6 Fourier modes and a spatial sparsity of 0.02. SAE-FNO consistently displays a broader correlation distribution with heavier tails compared to SAE-MLP.
Figure 19: FNO generalizes well out-of-distribution compared to SAE-MLP and SAE-CNN. Left: $p=1000$. Right: $p=256$.
Figure 20: Lifting accelerates learning. MNIST.
Figure 21: Lifting accelerates learning. CIFAR.
(a) $k=10$
(b) $k=25$
(c) $k=50$
(d) $k=100$
(e) $k=200$
(f) $k=500$
(g) $k=700$
(h) $k=1000$
Figure 22: Additional visualizations. SAE-MLP learned concepts on CIFAR.
(a) $k=10$
(b) $k=25$
(c) $k=50$
(d) $k=100$
(e) $k=200$
(f) $k=500$
(g) $k=700$
(h) $k=1000$
Figure 23: Additional visualizations SAE-FNO (modes = 6, spatially 1-sparse) learned concepts on CIFAR.
(a) 1-sparse
(b) 3-sparse
(c) 6-sparse
(d) 32-sparse
Figure 24: Additional visualizations. SAE-FNO (modes = 6) learned concepts on CIFAR as spatial sparsity changes.
Appendix C Appendix - Theoretical Analysis